Find the particular solution of ydx - 4dy = x ^ 2 dy when x = 4 and y = 1


1
Expert's answer
2022-03-02T12:40:07-0500
"ydx-4dy=x^2dy"

"(x^2+4)dy=ydx"

"\dfrac{dy}{y}=\dfrac{dx}{x^2+4}"

Integrate


"\int\dfrac{dy}{y}=\int\dfrac{dx}{x^2+4}"


"\ln|y|=\dfrac{1}{2}\tan^{-1}(x/2)+C"

When "x = 4" and "y = 1"


"\ln|1|=\dfrac{1}{2}\tan^{-1}(4/2)+C"

"C=-\dfrac{1}{2}\tan^{-1}(2)"

The particular solution is


"\ln|y|=\dfrac{1}{2}\tan^{-1}(x/2)-\dfrac{1}{2}\tan^{-1}(2)"